[Hallc_running] [New Logentry] RC update, and day shift is on standby

mack at jlab.org mack at jlab.org
Mon Jun 24 10:00:04 EDT 2019

Logentry Text:
Day shift was put on standby after doing the shift checklist. Hall C has gone to controlled access so you can potentially do useful, safe things. Magnets are still energized. Once we know the program, we can talk about what new angles and momenta to set. 

We're going to tread water for most of day shift since the ESR 15K supply still needs to be checked for contamination and a pressure check, and getting a high power target back is essential for us. If the ESR 15K is ready today, then we would also need to know if Hall A - which has cryo priority - will want to cool down first, delaying the start of the Hall C cooldown for a day. So our near-term future is uncertain due to two major decision branches, but in the worst case we'll have an alternative program  ready for swing shift with either no target or low power target. 

FYI, the draft alt program in prioritized order is

1) push for higher beam current now that Hall C is no longer merely opportunistic (Gaskell), 
2) if the operators run out of ideas on #1, and we're allowed to change the HMS angle, do an HMS pointing scan at 5 muA on carbon (Jones), 
3) sit at 5 muA on carbon for an extended period for further LAAPD tests (Jones, Xie). 

 There will be the usual RC meeting this afternoon at 4:30 pm in the Counting House. Perhaps by then we'll know which parallel universe we've landed in. Remote connection information is on the wiki



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3693007
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190624/862b8c33/attachment.html>

More information about the Hallc_running mailing list